### v1

# BulletGen: Improving 4D Reconstruction with Bullet-Time Generation

Abstract
Transforming casually captured, monocular videos into fully immersive dynamic experiences is a highly ill-posed task, and comes with significant challenges, e.g., reconstructing unseen regions, and dealing with the ambiguity in monocular depth estimation. In this work we introduce BulletGen, an approach that takes advantage of generative models to correct errors and complete missing information in a Gaussian-based dynamic scene representation. This is done by aligning the output of a diffusion-based video generation model with the 4D reconstruction at a single frozen “bullet-time” step. The generated frames are then used to supervise the optimization of the 4D Gaussian model. Our method seamlessly blends generative content with both static and dynamic scene components, achieving state-of-the-art results on both novel-view synthesis, and 2D/3D tracking tasks.
1 Introduction
The task of 3D reconstruction from color images is fundamental to computer vision, with a variety of solutions being proposed over the years [26, 51, 63, 64]. While impressive strides have been made towards improving both the completeness and accuracy of results, most methods and datasets focus exclusively on static scenes. However, the real world is dynamic, and the ability to reconstruct dynamic content opens up many new opportunities in immersive media generation and robotics.
The reconstruction of dynamic scenes presents a very challenging 4D problem, as it requires reasoning about both geometry and motion of the scene. This is exacerbated by the fact that traditional 4D reconstruction methods require capturing multi-view images of dynamic scenes with expensive camera rigs [53, 23], resulting in limited amount of useful datasets for algorithm development and evaluation. Therefore, building on developments in novel 3D scene representations [26, 51], the attention of the research community in recent years has focused on 4D reconstruction from single-view monocular videos [49, 54, 76]. But, as a monocular video only observes the scene from a single viewpoint at any given timestep, the problem of 4D reconstruction in this context is heavily under-constrained. As a consequence, existing methods can only find locally optimal solutions, and fail when synthesizing novel views substantially deviating from the training set views (Fig. 1).
Motivated by breakthroughs in learning priors from large-scale data, many recent methods have sought to mitigate the local minima problem by leveraging generative models – particularly diffusion-based methods [61, 67, 69] – to constrain the output to lie in the distribution of natural images. However, in the majority of cases, these methods only predict 2D projections of a dynamic scene, and therefore fail to take advantage of the efficient rendering and global consistency advantages provided by 3D scene representations that are optimized per-scene, such as Neural Radiance Fields (NeRFs) [51] and 3D Gaussians [26]. In this paper, we propose a method to incorporate these 2D projections into a consistent and plausible 4D reconstruction.
Our method, termed BulletGen, uses a video diffusion model to generate novel views for selected frozen timesteps – so-called bullet times [40, 72]. The diffusion model is conditioned on a rendered frame along with natural language image captions from the input video. The generated views are then used to iteratively supervise a global 3D representation. This latter step is achieved by accurately tracking and aligning the generated views to the input coordinate frame, and by using a novel reconstruction loss based on photometric, perceptual, semantic, and depth errors. By repeating this process for different time stamps, the initially inconsistent 2D projections are robustly incorporated into a consistent dynamic 3D reconstruction. We use 3D Gaussians as the underlying scene representation, but our method is general and could work with any differentiable 3D representation.
In summary, we make the following contributions:
  • We propose a method for dynamic 3D scene reconstruction from monocular RGB videos by augmenting the training views using a generative video diffusion model at selected bullet times.
  • We propose a robust reconstruction loss that allows us to align the generated views in a global coordinate frame, and use them to supervise a dynamic Gaussian reconstruction.
  • We compare our method against prior work, and show that BulletGen achieves state-of-the-art results on several benchmark datasets for novel view synthesis quality, and 2D/3D tracking accuracy.
  • The reconstructed dynamic scene also contains new synthesized parts of the scene that seamlessly and plausibly blend in the original scene for both the static (e.g. new pillows in the cat scene and generated walls in the skating scene in Fig. 1) and the dynamic parts (e.g. backside of the cat, the full head of the dog).
skating Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cat Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
dog Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Training view Our reconstruction SoM Our reconstruction SoM
Figure 1: Extreme novel view synthesis of a 4D scene with generative model guidance in frozen-time instances (bullet times). The input is only a monocular video on the left. We compare to the Shape-of-Motion (SoM) [73] method on the cat and dog sequences from the iPhone dataset [14] and the skating sequence from Nvidia dataset [87].
2 Related work
The problem of 3D reconstruction and novel-view synthesis has a long history in the fields of computer vision and graphics. Early approaches used dense multi-view images [8, 28, 75] to capture a spatio-angular light fields [18, 34], which could be re-parameterized to generate novel views via quadrilinear interpolation. However, the number of views required by sampling theory to enable such interpolation has been shown to be exorbitantly high [3]. As a result, novel view synthesis remained a niche task requiring expensive camera rigs, specialized sensors, or time-consuming capture setups [8, 16, 52, 75]. Mildenhall et al. [50] and Zhou et al. [95] were among the first to seek to overcome the sampling limits using learned data priors encoded as the weights of a convolutional neural network. Subsequent work built on this work by exploiting the ever-increasing capabilities of deep neural networks to not only lower the sampling requirements for view synthesis, even down to a single image  [4, 20, 27, 45, 47, 48, 68, 90], but also proposed new representations that could be optimized per scene to achieve exceptionally high quality and improved spatio-temporal consistency by exploiting scene-specific correlations across views  [49, 36, 30, 42]. These two research directions naturally converge on the task of 4D reconstruction from monocular videos. However, in this context, past work that has used per-scene optimization to track and correlate points across frames has often failed to take advantage of the strong data priors of pre-trained models specifically for view synthesis.
Per-scene monocular reconstruction. Following the seminal work of Mildenhall et al. [51] on static reconstruction by optimizing a scene-specific 3D representation by Neural Radiance Fields (NeRFs), many works [1, 13, 37, 54, 55, 59, 65, 66, 84] have adopted a similar strategy to exploit more global context for the problem of 4D reconstruction from a monocular video. The underlying representation often models the scene as an implicit field [79] parameterized by time, which is optimized using the input frames as supervision. The use of a single globally consistent representation, which also implicitly encodes a local smoothness prior, allows per-scene optimization to achieve impressive reconstruction results without leveraging any learned data priors [91]. However, as the optimization process can be poorly regularized and time-consuming, several methods have proposed using a monocular depth prior to guide the optimization process [9, 44, 62]. Furthermore, the introduction of 3D Gaussian splats [26] in recent years has led to methods that reconstruct either scene geometry [41, 76, 84, 44, 7], or both geometry and motion [36, 43, 49, 85, 31, 12], explicitly. Thus, a recent class of methods has additionally used 2D motion trajectories from pre-trained models to initialize and supervise the optimization of a scene-specific global representation [42, 33, 73]. However, these methods use learned priors primarily to regularize the optimization and to reason about the visible parts of the scene. As such, while they generate more stable reconstructions in these visible regions, they perform no better than previous methods on view extrapolation tasks.
Generative 4D reconstruction. Given the extremely high number of images required to meet the Nyquist sampling rate for view interpolation from multi-view images [3, 50], and the fundamentally under-constrained nature of the view extrapolation problem, many works have explored the use of learned data priors to circumvent these theoretical limitations [4, 50, 86, 90, 95]. In the extreme case, these methods aim for view synthesis from a single RGB image. The most effective prior for this task, in terms of output quality, turns out to be a learned conditional probability distribution over the space of natural images. This is represented as a deep neural network that generates samples from the underlying distribution [2, 17, 80]. Among this class of “generative” methods, Denoising Diffusion Probabilistic Models (DDPM) [21] have become especially popular due to their stability, mode coverage, quality, and flexibility [6]. Consequently, a number of recent works have used diffusion to set the bar of quality on 3D reconstruction and view synthesis from single or sparse images [88, 15, 46, 78, 70]. Subsequent work has extended the basic approach to 4D reconstruction from single images [67, 81, 94], sparse image sets [5, 94], and from monocular videos [61, 69, 89]. These methods rely on sampling all output frames from the generative model. Consequently, they do not provide accurate control over the camera movement and lack explicit spatio-temporal consistency constraints. In addition, the high memory requirements of video diffusion models limit these methods to short video sequences.
Concurrent work. CAT4D [77] and Vivid4D [22] are concurrent works with ours, and show the use of generative models to improve per-scene optimized monocular reconstructions. Both works proceed by first generating multi-view video from a monocular sequence, and then use it to supervise the optimization of a dynamic scene representation based on Gaussian splats. As such, the generation and optimization process is strongly decoupled. Our method uses an iterative approach where the generation steps alternate with the training of a Gaussian-based global 4D representation.
3 Method
Given $N$ frames of a monocular video sequence $\{I_{t}\}_{t=1}^{N}$ as input, our method acquires an initial 4D reconstruction of the scene as dynamic Gaussian splats. It then uses diffusion-based generative augmentation to obtain new observations for under-constrained regions, as well as synthesizing unseen parts of the scene. Using the original frames along with the generated views as supervision, the initial reconstruction is robustly optimized to obtain a photorealistic and globally consistent dynamic scene representation that allows for rendering of novel views at arbitrary time stamps and camera positions. Fig. 2 provides an overview of our pipeline.
Figure 2: BulletGen architecture. Starting from a monocular RGB video, we reconstruct the dynamic scene with Shape-of-Motion [73] given data-driven priors (motion masks, depths, long-term 2D tracks). Then, we generate novel views at selected frozen timesteps (bullet times) using a conditioned generative model. These generated views are localized and mapped to the current scene using an optimization based on photometric, perceptual, semantic, and depth errors. Final 4D reconstruction augments the scene and allows for higher quality extreme novel view synthesis.
3.1 Initial dynamic Gaussian Splatting reconstruction
To bootstrap the initial reconstruction, we build on top of several state-of-the-art methods. We choose dynamic Gaussian Splatting as a scene representation, which is an extension to the original 3D Gaussian Splatting (3DGS) [26] method. In particular, we use Shape-of-Motion [73], which relies on a range of scene priors extracted using methods like Track-Anything [30, 82] to extract masks for the moving objects, Depth Anything [83] to estimate relative monocular depth maps, UniDepth [58, 57] for metric depth alignment, and TAPIR [11, 10] for obtaining long-range 2D tracks of moving objects.
Each 3D Gaussian is represented by RGB color $c\in[0,1]^{3}$, its center position $\mu(t)\in\mathbb{R}^{3}$, rotation $r(t)\in\mathbb{R}^{4}$ (quaternion), scale $s\in\mathbb{R}^{3}$ and opacity $o\in[0,1]$. For static Gaussians, $\mu(t)$ and $r(t)$ are constant for all time stamps. For dynamic Gaussians, they are defined as a time-dependent weighted combination of a global set of motion bases [73] that are learned during optimization (usually a low number, e.g., 20). Standard 3DGS [26] can be used to render the dynamic Gaussians for a given time stamp $t$. Given the required camera pose, all Gaussians are sorted by distance from the camera center and then rendered by alpha-composing the splatted 2D projections, which are affine approximations of the exact 3D projection. Given the complexity and ill-posedness of the task due to limited observations, we do not model view-dependent illumination changes with spherical harmonics. In the initial bootstrapping phase, we run Shape-of-Motion for 1000 optimization epochs to obtain a dynamic scene reconstruction without any generative component (Fig. 2).
3.2 Generative augmentation
After initial scene reconstruction, we generate a set $\{G_{k}^{t}\}_{k=1}^{K}$ of $K$ novel views of the scene at selected times $t$, and select a conditioning view and target motion (explained below). We generate novel views of the dynamic scene at bullet time $t$, assuming that the scene is frozen in time. We use a controllable image-to-video diffusion model that is trained to create novel views of the scene given a single image and a text prompt, which is generated by Llama3 [19]. The conditioning prompt is a highly descriptive image captioning from the original monocular video, which serves as an anchor to be consistent with the original input. The used generative model also controls the motion by training different models for target motions, such as left and up directions. Then, the generated images $\{G_{k}^{t}\}_{k=1}^{K}$ are localized in the scene and used to update the scene with new observations.
Camera tracking. We estimate initial relative camera poses between frames of the same bullet time using the recently introduced VGGT [71] transformer. Since VGGT-predicted depth is inaccurate, we estimate state-of-the-art monocular depth from MoGe [74] and align it to VGGT depth. As MoGe/VGGT depths are not metric, we estimate a single scaling factor to align them to the current 4D reconstruction by minimizing the depth and RGB reprojection error in the covisible regions, producing final depth estimates $\{D_{k}\}_{k=1}^{K}$. At this stage, the camera poses of generated views are roughly aligned to the current scene reconstruction, but the alignment is not yet pixel-perfect, which is crucial for the following scene augmentation. To achieve pixel-perfect alignment, we leverage the Gaussian Splatting SLAM method SplaTAM [25] for accurate camera tracking. Camera intrinsics for the generated cameras are assumed to be equal to the estimated intrinsics of the original monocular video. Extrinsics $\mathrm{\textbf{E}}_{k}\in\mathbb{R}^{4\times 4}$ for generated views $G_{k}^{t}$ are initialized to the scale-aligned VGGT estimates. Then, the scene is rendered using the rendering function $\mathcal{R}(\mathrm{\textbf{E}}_{k})$ producing RGB images $\mathcal{R}_{I}(\mathrm{\textbf{E}}_{k})$ and depth images $\mathcal{R}_{D}(\mathrm{\textbf{E}}_{k})$. Additionally, we render silhouettes $\mathcal{R}_{S}(\mathrm{\textbf{E}}_{k})$ by accumulating Gaussian densities along the ray, which are used to define visibility masks as $V_{k}=\mathcal{R}_{S}(\mathrm{\textbf{E}}_{k})>0.99$.
Robust loss and optimization. The minimized loss function consists of four terms representing the photometric, perceptual, semantic, and depth errors. The photometric loss is the L1 loss in the RGB space, the perceptual loss is the LPIPS loss [92] based on AlexNet [32] features, and the semantic loss is the cosine similarity between CLIP scores [60] of the renderings $\mathcal{R}_{I}(\mathrm{\textbf{E}}_{k})$ and the generated images $G_{k}$. Finally, the depth loss is the L1 distance between depth renderings and the aligned depth maps. Since the generated images are usually not perfectly 3D consistent in the pixel space, we put the highest weight on the semantic and perceptual losses and keep the photometric loss only as a low-level learning signal. The final loss function is the weighted sum:
($\begin{split}\mathcal{L}(\{\mathrm{\textbf{E}}_{k}\}|V_{k})=\sum_{k=1}^{K}&\alpha_{1}\text{L1}\Bigl{(}G_{k},\mathcal{R}_{I}(\mathrm{\textbf{E}}_{k})\Big{|}V_{k}\Bigr{)}+\alpha_{2}\text{LPIPS}\Bigl{(}G_{k},\mathcal{R}_{I}(\mathrm{\textbf{E}}_{k})\Big{|}V_{k}\Bigr{)}\\ +&\alpha_{3}\text{CLIP}\Bigl{(}G_{k},\mathcal{R}_{I}(\mathrm{\textbf{E}}_{k})\Big{|}V_{k}\Bigr{)}+\alpha_{4}\text{L1}\Bigl{(}D_{k},\mathcal{R}_{D}(\mathrm{\textbf{E}}_{k})\Big{|}V_{k}\Bigr{)}\,,\end{split}$ (1)
where we compute the loss only over the visible area $V_{k}$. We optimize camera poses $\{\mathrm{\textbf{E}}_{k}\}$ by minimizing the loss for 100 epochs, starting from the initialization described above. All other parameters, including the scene representation, are fixed. Once the camera poses are optimized, producing $\{\mathrm{\textbf{E}}_{k}^{*}\}$, it mostly leads to sufficiently accurate, pixel-wise alignment. To remove bad estimates, we keep only the first $K^{\prime}$ generated views up until the first generative view whose loss $\mathcal{L}$ (1) is above a threshold $\gamma$, which is set in a conservative way so only well-aligned views are retained. After this step, we have $K^{\prime}\leq K$ generated views $\{G_{k}\}$, their depths $\{D_{k}\}$, and optimized camera poses $\{\mathrm{\textbf{E}}_{k}^{*}\}$.
right view frontal view left view right view frontal view left view
depth at 40/180 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
time 40/180 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
time 80/180 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
time 160/180 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Inputs Shape-of-Motion Ours
Figure 3: Extreme novel view synthesis across space and time. The generation for training was performed $n_{G}=7$ times for $n_{S}=5$ bullet-time stamps, i.e. 1, 45, 90, 135, 180 for a sequence with length of 180 frames. The renderings shown here are at time stamps and viewpoints that were not generated using a generative model, which shows that using only several bullet-time reconstructions is enough to reconstruct a dynamic scene. Temporal slices for all 180 time stamps are shown in Fig. 5.
Densification. After obtaining well-aligned generative views, we now proceed to updating the scene. First, we perform Gaussian densification based on the densification mask proposed in SplaTAM [25]:
$M_{k}=\Bigl{(}\mathcal{R}_{S}(E_{k}^{*})<0.5\Bigr{)}+\Bigl{(}D_{k}<\mathcal{R}_{D}(E_{k}^{*})\Bigr{)}\Bigl{(}L_{1}(D_{k},\mathcal{R}_{D}(E_{k}^{*}))<\lambda\text{MDE}\Bigr{)}\,,$ (2)
where we densify areas with insufficient density or where new geometry is in front of the current geometry, unless it is due to noise as measured by $\lambda=50$ times the median depth error [25]. This way, we initialize a new Gaussian at every pixel in the densification mask based on its color and depth. For the newly densified Gaussians, we decide if they are static or dynamic based on the nearest neighboring Gaussian’s static/dynamic label, where the dynamic Gaussians’ locations are moved to the currently generated bullet time stamp. The weights for the motion bases of new dynamic Gaussians are initialized to the nearest neighbor.
Scene update. To update the reconstructed scene, we optimize a weighted joint loss, which is a sum of the proposed tracking loss $\mathcal{L}$ (1) on the generated views, including all previous generated views over all bullet time stamps, and the SoM loss function to always keep the scene consistent with the original video. Since the scene is already densified, we do not need the visibility masks $V_{k}$ in this stage and, thus, the loss is computed over the entire generated image, as compared to the loss used for camera tracking. We optimize this joint loss function for 100 epochs by optimizing parameters of all Gaussians and the motion bases. After this, we repeat it all with new generated views (Fig. 2).
Time selection. One of the hyperparameters of our method is the number $n_{S}$ of sampled bullet time stamps. We sample them uniformly between the first and last time stamps, i.e. $t=1$ and $t=N$. In order to guarantee smooth motion, it is important to keep some distance between sampled time stamps since each bullet time generation may contain slight inconsistencies. If the distance between two sampled time stamps is low, this can generate effects similar to jittering. However, if the distance is sufficiently large, it gives enough time for smooth transitions between two bullet-time reconstructions with their own slight inconsistencies. For robust estimation, we start with the middle frame bullet-time stamp, e.g. $N/2$. Then, we proceed with the furthest time stamps. For instance, when $N=100$ and $n_{S}=5$, the sampled time stamps are $\{1,25,50,75,100\}$, but in the following order $\{50,1,100,25,75\}$. Also note that the bullet-time stamp selection can be progressive and proceed auto-regressively, and in our example, can continue with additional time stamps $\{12,37,62,87\}$ for $n_{S}=9$.
View selection. The used generative model currently supports only static scenes and comes with two operating modes for generating left- and upward trajectories. By flipping the image horizontally, applying the leftward model, and flipping back, we can also generate rightward trajectories. Note that we do not apply the same principle to the upward direction, because the model only works well with upright imagery. Thus, we select between three modes for the view selection, i.e. left, right, up. Once bullet time is selected, we proceed with $n_{G}$ generations for the selected time. We tested three modes. First, for $n_{G}=3$, we select {up, left, right} in this order. For $n_{G}=5$, we select {up, left, right, left, right}, and for $n_{G}=7$, the used sequence is {up, left, up, right, up, left, right}. The view selection for generative model conditioning for each chosen direction is as follows. For leftward motion, we choose the left-most image (either the original one or the generated one), where the angle is measured from the middle of the original camera poses and viewing at the middle of the dynamic scene. Similarly, for rightward motion, we choose the right-most image. In all cases, only images at the currently selected bullet-time stamp are considered, either generated or the original one (there is only one original image for every time stamp). For upward motion, we choose the last generated image at this bullet time (or the original one if this is the first generation).
SoM Ours GT SoM Ours GT
spin Refer to caption Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption
paper-windmill Refer to caption Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption
balloon2 Refer to caption Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption    Refer to caption   Refer to caption   Refer to caption
Training view Testing view 1 Testing view 2
Figure 4: Qualitative evaluation on several benchmark datasets. The input is a monocular video as shown in the training view column. Both benchmark datasets (Nvidia [87] and DyCheck iPhone [14]) have several additional testing cameras. We compare novel view synthesis to Shape-of-Motion (SoM) [73] and zoom-in to highlight the differences. Our method is able to provide more accurate and sharper reconstructions, both in static and dynamic parts of the scene.
Implementation details. All optimizations are performed using the ADAM optimizer [29]. The batch size for the generative views and the original video frames is set to 8 (thus, we have two concatenated batches of 8 in every iteration). The hyperparameters in Eq. (1) have been empirically determined and fixed to $\alpha_{1}=0.02$, $\alpha_{2}=0.1$, $\alpha_{3}=0.1$, $\alpha_{4}=0.5$ for all experiments. The threshold for generative views is set to $\gamma=0.4$. For the SoM loss function on the original video, we use their default weights. The total number of generated views per generation is fixed at $K=50$. The number of generations used is $n_{G}=7$, while the number of selected bullet times is $n_{S}=9$, unless specified otherwise. The average optimization time on a sequence from the iPhone dataset [14] with full resolution takes around 3 hours (including initial SoM optimization of 1.5 hours) on an Nvidia A100 80GB GPU. All main experiments required around 1 week of A100 GPU time, and preliminary non-included experiments required an additional month of GPU time. For evaluation on custom data, we use MegaSaM [38] to estimate camera poses. For evaluation on benchmark datasets, we used provided poses computed from COLMAP [63, 64] for fair comparison to other methods.
4 Experiments
We evaluate on the DyCheck iPhone dataset [14] and Nvidia dynamic dataset [87], which are commonly used datasets for novel view synthesis evaluation in dynamic scenes.
View synthesis metrics. We evaluate the standard PSNR, SSIM, and LPIPS scores between the ground truth and the rendered novel views. On the iPhone dataset, we measure those metrics on the covisible regions for fair comparison to other methods. As mentioned by [42], these scores are not robust enough on the current datasets due to camera misalignment, color differences between the training and test cameras, as well as the nature of the task. Thus, as proposed in [42], we also report CLIP-I [60] scores, which we see as more representative in this ill-posed task.
Tracking metrics. The iPhone dataset contains long-range 3D point tracking annotations. Thus, we also measure 3D end-point-error (EPE) at every timestep and the percentage of points that are within a certain radius of the ground truth 3D point (denoted by $\delta_{3D}^{\tau}$, where $\tau$ is either 5cm or 10cm). By projecting those 3D tracking annotations, we obtain 2D tracking annotations and report Average Jaccard (AJ) metric, average position accuracy ($<\delta_{avg}$) and Occlusion Accuracy (OA) metric.
Baselines. We compare to a range of methods that are specifically designed for dynamic scene reconstruction. Methods based on Neural Radiance Fields (NeRF) [51] and its variants are T-NeRF [35], HyperNeRF [56], and DynIBaR [39]. More recently, methods based on Gaussian Splatting were proposed, such as Deformable-3DGS [84], 4D-GS [76], Shape-of-Motion [73], and HiMoR [42]. We also compare to methods based on generative diffusion models, such as CoCoCo [96], StereoCrafter [93], ViewCrafter [88], CAT4D [77], and Vivid4D [22]. Many of these methods are concurrent with our work, are only pre-prints (not published), and do not provide code, e.g. [77, 22, 42]. Nevertheless, we strive to provide insightful comparisons by evaluating on the reported datasets chosen by these methods. For 2D and 3D tracking scores, we additionally compare to CoTracker [24] and TAPIR [11], which are lifted to 3D by DepthAnything [83].
3D Tracking 2D Tracking
[…]
HyperNeRF [56] 0.182 28.4 45.8 10.1 19.3 52.0
DynIBaR [39] 0.252 11.4 24.6 5.4 8.7 37.7
Deformable-3D-GS [84] 0.151 33.4 55.3 14.0 20.9 63.9
CoTracker [24] + DA [83] 0.202 34.3 57.9 24.1 33.9 73.0
TAPIR [11] + DA [83] 0.114 38.1 63.2 27.8 41.5 67.4
Shape-of-Motion [73] 0.082 43.0 73.3 34.4 47.0 86.6
Ours 0.071 51.6 77.6 36.6 49.5 87.4
Table 1: Evaluation on the iPhone dataset, 3D and 2D tracking. Our method outperforms all state-of-the-art methods in terms of 2D and 3D tracking accuracy as measured by various metrics.
[…]
T-NeRF [35] 15.60 0.55 0.55 0.86
HyperNeRF [56] 15.99 0.59 0.51 0.87
Deform.-3DGS [84] 11.92 0.49 0.66 0.79
SoM [73] 16.72 0.63 0.45 0.86
HiMoR [42] (no code) - - 0.46 0.89
CAT4D [77] (no code) 17.39 0.61 0.34 -
Ours 16.78 0.64 0.39 0.90
Table 2: Evaluation on the iPhone dataset, novel view synthesis. The proposed method achieves state-of-the-art performance in terms of novel view synthesis.
[…]
4D-GS [76] 14.01 0.39 0.59
CoCoCo [96] 14.99 0.47 0.53
StereoCrafter [93] 14.85 0.49 0.57
ViewCrafter [88] 14.94 0.49 0.58
SoM [73] 14.56 0.46 0.53
Vivid4D [22] 15.20 0.50 0.49
Ours 16.38 0.51 0.45
Table 3: Evaluation on the iPhone dataset, subset chosen by Vivid4d [22]. Our method outperforms all other methods.
The iPhone dataset. The iPhone dataset [14] contains 12 sequences with a moving training camera captured with hand-held iPhone device, two static test cameras (for 5 sequences), and 3D tracking annotations. As shown in Table 1, our method outperforms all other state-of-the-art methods on all metrics in terms of 2D and 3D tracking scores. This is achieved by providing more constraints for scene geometry, which allows for better 3D triangulation during scene reconstruction. We also achieve state-of-the-art scores in terms of novel view synthesis (Table 3). For the sake of completeness, we add results reported in CAT4D [77] (unpublished work), which achieves slightly better PSNR and LPIPS scores, whereas our method obtains better SSIM results. However, we want to stress that the table remains incomplete in terms of CLIP-I score due to a lack of publicly accessible code and missing information about the training procedure and its overall training scale. To compare to the recently released Vivid4D [22] method (also without publicly available code), we choose their subset of the iPhone dataset that was used in their evaluation. Our proposed method again outperforms other methods (Table 3).
[…]
SoM 15.26 0.454 0.388 0.87
Ours 17.02 0.462 0.386 0.87
Table 4: Evaluation on the Nvidia dataset. Our method improves the Shape-of-Motion (SoM) reconstruction on all metrics.
The Nvidia dataset. The Nvidia dataset contains 9 dynamic scenes captured by 12 synchronized static cameras. We use camera 1 as input, and cameras 2 and 3 as output for evaluation. This dataset contains testing views close to the training views, therefore the CLIP-I and LPIPS metrics do not show significant improvement, but in terms of PSNR, our improvement w.r.t. Shape-of-Motion is significant (see Table 4).
[…]
SoM 16.72 0.63 0.45 0.86
number of generations $n_{s}=9$, $n_{g}=3$ 16.81 0.63 0.40 0.89
$n_{s}=9$, $n_{g}=5$ 16.80 0.63 0.40 0.89
$n_{s}=9$, $n_{g}=7$ 16.78 0.64 0.39 0.90
number of bullet-times $n_{s}=3$, $n_{g}=7$ 15.95 0.62 0.45 0.88
$n_{s}=5$, $n_{g}=7$ 16.36 0.62 0.44 0.89
$n_{s}=9$, $n_{g}=7$ 16.78 0.64 0.39 0.90
Table 5: Ablation study (iPhone dataset). The number of generations $n_{G}$ per time stamp slightly improves most metrics, and the number of selected bullet-time stamps $n_{S}$ improves all metrics.
Ablation study. The main hyperparameters of our method are the number of generations $n_{G}$ per selected bullet-time stamp, and the total number of those time stamps $n_{S}$. Table 5 shows that increasing $n_{G}$ per time stamp slightly improves most metrics, but is not significant given that the test cameras are not far from the training ones. A high value for $n_{G}$ is important for extreme novel view point synthesis. However, the number of selected bullet time stamps $n_{S}$ is vital even in the standard setting, and a higher number of time stamps improves all metrics.
Qualitative results. Qualitative results for the iPhone dataset are provided in Fig. 1 (cat, dog) and Fig. 4 (spin, paper-windmill), and for the Nvidia dataset in Fig. 1 (skating) and Fig. 4 (balloon2). As can be seen in Fig. 4, the testing views in the benchmark datasets are close to the training views. However, our method allows for much more extreme novel view synthesis, going beyond what is possible by other methods. Unfortunately, there is no dataset with such extreme view point testing cameras, which can be seen as a limitation in this emerging field. Fig. 3 shows qualitative examples of extreme novel views. BulletGen significantly improves rendering quality in such cases. Temporal slices in the $x$ direction are shown in Fig. 5, which highlights that our method produces smooth and consistent rendering results over time, especially when compared to SoM.
Figure 5: Temporal plane slices of extreme camera views. Visualizing the highlighted rows across time in $xt$ space shows that our method suffers from fewer temporal artifacts than Shape-of-Motion (SoM) caused by floating Gaussians and exploding geometry. Our rendered view is shown on the left.
Limitations. BulletGen assumes that the initial Shape-of-Motion optimization performs reasonably well, at least from the viewpoints of the original video, which in turn depends on how well all monocular priors were estimated.
5 Conclusion
We introduced a method for dynamic 3D reconstruction from a monocular video, effectively addressing extreme novel view synthesis challenges with drastic improvements over the state of the art. By integrating static bullet-time video generation with dynamic 3D Gaussian splatting our approach enhances both 2D and 3D tracking accuracy, and novel view synthesis. Our method seamlessly incorporates new scene elements, improving rendering quality. Our results on standard benchmark datasets confirm the efficacy of this approach for dynamic scene reconstruction in complex environments.